Self Check? Solve: _aa_ 225 + 34a2-0. (Enter your answers from smallest to largest.)
Answer #1

Answers: (From smallest to the larget values)

a=-5  

a=-3

a=+3

a=+5

Explanation: The given equation is

-a^4-225+34a^2=0

multiplying by - sign on both sides

a^4+225-34a^2=0 ------------------------------------- (1)

Now, let us substitute a2=x . so, a4=x2. So, equaation 1 becomes

x^2+225-34x=0

x^2-34x+225=0

This equation can be solved by factoring the middle term. There are two numbers -9 and -25 such that (-9)*(-25)=225 (product of first and last term) and -9-25=-34 (middle term). So,

x^2-25x-9x+225=0

x(x-25)-9(x-25)=0  

(x-25)(x-9)=0 ---------------------------------------------- (2)

So, we have

(x-25)=0\Rightarrow x-25+25=0+25\Rightarrow x=25 ---------------------------- (3)

(x-9)=0\Rightarrow x-9+9=0+9\Rightarrow x=9 ---------------------------------- (4)

Now, since a2=x. So,

(1) using x=25 in a2=x

a^2=25

taking square root on both sides

\sqrt{a^2}=\sqrt{25}

a=\pm 5

So, a=+5 and a=-5

(2) Using x=9 in a2=x

a^2=9

taking square root on both sides

\sqrt{a^2}=\sqrt{9}

a=\pm3   

So, a=+3 and a=-3

For checking substiture the values of a=+3, -3, +5 and -5 one by one and check the validity of eqn 1.